In a telegraph cable, the measure of the speed of the signal is proportional to π‘₯^2 𝑙𝑛 (1/π‘₯), where

π‘₯ is the ratio of the measure of the radius of the core of the cable to the measure of the thickness of the cable’s winding. Find the value of 𝑙𝑛π‘₯ for which the speed of the signal is greatest.


1
Expert's answer
2020-12-17T18:45:38-0500

I presume the function is:


t(x)=x2ln(x). You are looking for a maximum of t(x). This will happen if t'(x) = 0. Thus calculate the first derivative. Use the product rule:


t'(x)=ln(x)β‹…2x+x21/x=x(2ln(x)+1)

Now t'(x) =0. Solve for x.

x(2ln(x)+1)=0.

x=0 this is not allowed or ln(x)=-1/2.

The answer is: ln(x)=-1/2